Grade 11General Physics

If a particle has initial velocity of 3i+4j and an acceleration of 0.3i +0.4j it`s speed after 10 s is ???

Don
8 Years agoGrade 11

1 Answer

IHate Ppts
ApprovedApproved Tutor Answer8 Years ago
Vx=Ux+AxT
V=3+0.3(10)
Vx=6 m/s
Vy=4+0.4(10)
Vy=8 m/s
Answer is 6i+8j
Or 10 m/s.
Hope this helped. Cheers!